solve by substitution method
6x+7y= -13
-4x+y=37

Answer by user_dude2008(1862)

-4x+y=37
y=4x+37


6x+7y= -13
6x+7(4x+37)= -13
6x+28x+259=-13
34x+259=-13
34x=-13-259
34x=-272
x=-8

y=4x+37=4(-8)+37=-32+37=5

Answers: x=-8, y=5